Consider the case of Alexander Industries:

Alexander Industries is considering a project that requires an investment in new equipment of $3,570,000. Under the new tax law, the equipment is eligible for 100% bonus depreciation at t = 0 so the equipment will be fully depreciated at the time of purchase. Alexander estimates that its accounts receivable and inventories need to increase by $680,000 to support the new project, some of which is financed by a $272,000 increase in spontaneous liabilities (accounts payable and accruals). The company's tax rate is 25%.

The after-tax cost of Alexander’s new equipment is   .

Alexander’s initial net investment outlay is   .

Suppose Alexander’s new equipment is expected to sell for $400,000 at the end of its four-year useful life, and at the same time, the firm expects to recover all of its net operating working capital (NOWC) investment. Remember, that under the new tax law, this equipment was fully depreciated at t = 0. If the firm’s tax rate is 25%, what is the project’s total termination cash flow?

$400,000

$508,000

$300,000

$708,000
Answer #1

After tax cost of Equipment = Purchase cost – Tax savings on Depreciation

= 3,570,000 – 3,570,000*25%

= $2,677,500

Initial Outlay = Cost of Equipment + Increase in current assets – Increase in current liabilities

= 2,677,500 + 680,000-272000

= $3,085,500

Termination Cash flows:

After tax salvage value = (400,000-0)*(1-25%) = $300,000

Recovery of Working Capital = $408,000

Termination Value = $708,000